OpenStudy (anonymous):

Find S11 for the geometric series 1 + 3 + 9 + 27 +… A. 84,934 B. 88,573 C. 93,198 D. 96,143

OpenStudy (texaschic101):

sn = a1 (1 - r^n) / (1 - r) n = sum of n terms = 11 a1 = first term = 1 r = common ratio = 3 now we sub s11 = 1(1 - 3^11) / (1 - 3) s11 = 1 (1 - 177147) / -2 s11 = 1 (-177146) / -2 s11 = -177146/-2 s11 = 88573